# Balancing Embedding Spectrum for Recommendation

### v1

@@
Abstract
Modern recommender systems heavily rely on high-quality representations learned from high-dimensional sparse data. While significant efforts have been invested in designing powerful algorithms for extracting user preferences, the factors contributing to good representations have remained relatively unexplored. In this work, we shed light on an issue in the existing pair-wise learning paradigm (i.e., the embedding collapse problem), that the representations tend to span a subspace of the whole embedding space, leading to a suboptimal solution and reducing the model capacity. Specifically, optimization on observed interactions is equivalent to a low pass filter causing users/items to have the same representations and resulting in a complete collapse. While negative sampling acts as a high pass filter to alleviate the collapse by balancing the embedding spectrum, its effectiveness is only limited to certain losses, which still leads to an incomplete collapse. To tackle this issue, we propose a novel method called DirectSpec, acting as a reliable all pass filter to balance the spectrum distribution of the embeddings during training, ensuring that users/items effectively span the entire embedding space. Additionally, we provide a thorough analysis of DirectSpec from a decorrelation perspective and propose an enhanced variant, DirectSpec+, which employs self-paced gradients to optimize irrelevant samples more effectively. Moreover, we establish a close connection between DirectSpec+ and uniformity, demonstrating that contrastive learning (CL) can alleviate the collapse issue by indirectly balancing the spectrum. Finally, we implement DirectSpec and DirectSpec+ on two popular recommender models: MF and LightGCN. Our experimental results demonstrate its effectiveness and efficiency over competitive baselines.
Index Terms:
@@ Abstract
I Introduction
@@
Recommender systems have penetrated into our daily life, we can see them everywhere such as e-commerce [1], short-video [2], social network [3], and so on. Collaborative filtering (CF), a fundamental technique for recommendation to discover user preference based on the historical data, has attracted much attention in the last decades. The most extensively used CF technique, matrix factorization (MF) [4] which represents users and items as low dimensional latent vectors and estimates the rating as the inner product between latent vectors, has been the cornerstone of modern recommender systems. Since MF estimates the rating with a simple linear function, subsequent works mostly focus on designing more powerful and complex algorithms to model non-linear user-item relations, including but not limited to multi-layer perceptron (MLP) [5], attention mechanism [6], reinforcement learning [7], transformer [8], diffusion model [9] graph neural network (GNN) [10, 11], etc., and have shown tremendous success.
@@
Although different kinds of methods have been proposed, most of them can be considered as MF variants whose goal is to learn low dimensional representations (with dimension $d$) from the high dimensional sparse interaction matrix (with dimension $D\gg d$). Figure 1 illustrates the top 500 normalized singular value distribution of the interaction matrix of CiteULike (see Section 5.1 for data description). We observe that users/items are predominantly distributed along a few dimensions in the original space while most dimensions barely contribute (i.e., with singular values close to 0) to the representations. Thus, when users and items are mapped into a more compact embedding space, it is expected that redundant dimensions are all removed and each dimension contributes to the user/item representations as equally and uniformly as possible (i.e., the representations make full use of the embedding space).
@@
Existing recommendation methods are mostly optimized by pulling observed user-item pairs closer than unobserved ones. Here, a simple yet fundamental question arises: Can user/item representations of existing work make full use of all dimensions? Unfortunately, by analyzing the spectrum of the embedding matrix, we empirically and theoretically show that users/items tend to span a subspace of the whole embedding space (with dimension $d^{\prime}<d$), where the embeddings collapse along all (complete collapse) or certain dimensions (incomplete collapse). Particularly, optimization solely on observed interactions is equivalent to a low pass filter, where the representations of users and items tend to collapse to a constant vector. Negative sampling is the most common technique to optimize recommendation algorithms without causing an explicit embedding collapse by pushing away the unobserved user-item pairs, and we show that it is equivalent to a high pass filter that decelerates the speed of collapse by balancing the embedding spectrum. However, the effectiveness of negative sampling is only limited to certain loss functions such as Bayesian personalized ranking (BPR) [12] and binary cross-entropy (BCE) loss [5], where collapse over certain dimensions still exists. Due to the data sparsity and long tailed distribution, increasing negative sampling ratios has been shown effective improving the representation quality [5, 13], whereas we also empirically demonstrate that it cannot further alleviate the collapse issue by evaluating different negative sampling ratios.
@@
Figure 1: An illustration of the data distribution in the original space and expected distribution in the learning embedding space.
@@
In this work, we tackle the embedding collapse issue from a spectral perspective. We observe that the extent of the collapse is closely related to the spectrum distribution of the embedding matrix. Specifically, only one singular value dominates when the representations completely collapse, whereas the singular values are uniformly distributed when the representations make full use of the embedding space. Inspired by this observation, we propose a novel method dubbed DirectSpec acting as an all pass filter to ensure that all dimensions equally contribute to the representations. We theoretically and empirically show that DirectSpec can completely prevent the embedding collapse without explicitly sampling negative pairs by directly balancing the spectrum distribution, and provide a simple implementation with a complexity only as $\mathcal{O}(B^{2}d)$ where $B$ is the batch size. Moreover, we shed light on DirectSpec from a decorrelation perspective, and propose an enhanced variant DirectSpec+ which employs self-paced gradients to optimize the irrelevant samples that are highly correlated more effectively. By showing the close connection between DirectSpec+ and uniformity, we discover that contrastive learning (CL) can alleviate embedding collapse by balancing spectrum distribution in a similar way to DirectSpec, explaining the effectiveness of CL based recommendation algorithms. Finally, we implement DirectSpec and DirectSpec+ on two popular baselines: MF [4] and LightGCN [14]. Experimental results show that DirectSpec+ improves BPR and LightGCN by up to 52.6% and 41.8% in terms of nDCG@10, respectively. The contribution of this work can be summarized as follows:
@@
  • We theoretically and empirically show that existing recommendation methods suffer from embedding collapse, that the representations tend to fall into a subspace of the whole embedding space, and analyze the causes behind this issue.
@@
  • We propose a novel method dubbed DirectSpec which acts as an all pass filter to directly balance the embedding spectrum. We empirically and theoretically show that DirectSpec can prevent embedding collapse.
  • We unveil the effectiveness of CL by showing that uniformity, a key design of CL can alleviate collapse issue by indirectly balancing the spectrum distribution that can be considered as a special case of DirectSpec.
@@
  • Extensive results on three datasets demonstrate the efficiency and effectiveness of our proposed methods.
II Preliminaries
II-A Pairwise Learning for CF
@@
Given the user set $\mathcal{U}$ and item set $\mathcal{I}$, the interaction matrix is defined as $\mathbf{R}\in\{0,1\}^{\left|\mathcal{U}\right|\times\left|\mathcal{I}\right|}$, the observed interactions are represented as $\mathbf{R}^{+}=\{r_{ui}=1|u\in\mathcal{U},i\in\mathcal{I}\}$. Users and items are first mapped to low dimensional vectors $\mathbf{H}\in\mathbb{R}^{(\left|\mathcal{U}\right|+\left|\mathcal{I}\right|)\times d}$ through an encoder, where the encoder can be as simple as a linear mapping [4] or advanced algorithms such as MLPs [5], GNNs [15], etc. Let $\mathbf{h}_{u}$ and $\mathbf{h}_{i}$ be the $u$’s and $i$’s representations, respectively. The goal of CF is to predict unobserved interactions $\mathbf{R}^{-}=\{r_{ui}=0|u\in\mathcal{U},i\in\mathcal{I}\}$ estimated as the inner product between the user and item representations: $\hat{r}_{ui}=\mathbf{h}_{u}^{T}\mathbf{h}_{i}$. The model parameters $\Theta$ are optimized through a loss function $\mathcal{L}$ formulated as follows:
${\rm arg}\mathop{{\rm min}}\limits_{\mathbf{\Theta}}\mathcal{L}(\hat{r}_{ui},r_{ui}).$ (1)
@@
The loss function measures the difference between the estimated score and the ground truth. BPR and BCE loss are two extensively used learning frameworks for CF methods:
$\mathcal{L}_{BPR}=\sum_{(u,i)\in{\mathbf{R}}^{+},(u,j)\in{\mathbf{R}}^{-}}-\ln\sigma(\hat{r}_{ui}-\hat{r}_{uj}),$ (2)
@@
where $\sigma(\cdot)$ is the sigmoid function. BPR loss maximizes the difference between observed and unobserved user-item pairs, while BCE loss directly pulls the observed pairs close and pushes the unobserved ones away from each other. The embeddings are a low dimensional approximation of the sparse high dimensional interaction matrix, thus they should contain diverse and rich information representing the user-item relations. Rank is a commonly used metric to measure the dimension of a matrix, while it fails to tell the difference between the embedding matrix (1) with a ‘sphere’ distribution that users/items are uniformly distributed in each dimension of the space and (2) with a ‘spheroid’ distribution that users/items are predominantly distributed over some dimensions and insignificantly distributed over other dimensions. Although both matrices have the same rank, apparently (1) contains more diverse information than (2). Here, we introduce another tool:
@@
$erank(\mathbf{H})=exp\left(H\left(p_{1},\cdots,p_{d}\right)\right),$ (3)
@@
Compared with rank, erank takes the singular value distribution into consideration: the more uniform (sharper) of the distribution, the higher (lower) of the erank [16]. The embedding matrix contains the least information when there is only one leading non-zero singular values ($\sigma_{2}=\cdots=\sigma_{d}=0$, and $erank(\mathbf{H})=rank(\mathbf{H})=1$), while erank is maximized when each dimension equally contributes to the representations: $\sigma_{1}=\cdots=\sigma_{d}$ ($erank(\mathbf{H})=rank(\mathbf{H})=d$).
II-B Graph Representation and Graph Filtering
@@
Graph Filtering. Let $11=\lambda_{1}>\lambda_{2}>\cdots>\lambda_{n}=-$ 1 be the eigenvalues of $\mathbf{A}$ where $n=|\mathcal{V}|$. The component closer to $\lambda_{1}$ and $\lambda_{n}$ correspond to a lower and higher frequency, respectively. $\mathcal{F}^{L}(\mathbf{A})$ is a low pass filter if $|\mathcal{F}^{L}(\lambda_{i})|>|\mathcal{F}^{L}(\lambda_{j})|$ for $\lambda_{i}>\lambda_{j}$ and $\mathcal{F}^{H}(\mathbf{A})$ is a high pass filter if $|\mathcal{F}^{H}(\lambda_{i})|>|\mathcal{F}^{H}(\lambda_{j})|$ for $\lambda_{i}<\lambda_{j}$.
Note that $\mathcal{F}^{L}(\mathbf{A})$ is defined as a low pass filter as it emphasizes more on the lower frequencies and similarly for the definition for high pass filters. Please refer to [17] for detailed introduction of graph frequency.
Figure 2: Embedding collapse on Yelp ($d=64$). (a) and (b) show how embeddings completely and incompletely collapse, respectively; (c) and (d) show how nDCG@10 changes as the training proceeds under the two situations.
III Embedding Collapse in CF
III-A Complete Collapse
Let us first consider a extensively used loss function in recommendation: log loss. Here, we only optimizes the observed interactions: $\mathcal{L}=-\sum\ln\sigma(\hat{r}_{ui}),(u,i)\in\mathbf{R}^{+}$.
@@ III-A Complete Collapse
$\mathbf{H}^{(l+1)}=\mathbf{H}^{(l)}-\alpha\frac{\partial\mathcal{L}}{\partial\mathbf{H}^{(l)}}=\mathbf{H}^{(l)}+\alpha\mathbf{A}\mathbf{H}^{(l)},$ (4)
@@
where $\alpha\in(0,1)$, $\mathbf{A}_{ui}=1-\sigma(\hat{r}_{ui})$ for $(u,i)\in\mathbf{R}^{+}$. Equation (4) is similar to the message passing in GNN [18], which makes the nodes similar to their neighborhood. Repeating Equation (4) multiple times can further reach the higher-order neighborhood, causing indirectly connected nodes to be similar. While $\mathbf{A}_{ui}$ controlling the magnitude of the gradient tends to vanish as training proceeds, and eventually $\mathcal{L}$ converges when $\mathbf{A}$ degenerates into a zero matrix. Note that this issue does not exist on other loss functions such as Euclidean distance loss: $\sum\left\|\mathbf{h}_{u}-\mathbf{h}_{i}\right\|^{2}_{2}$ with fixed magnitude of gradients, where Equation (4) can be rewritten as:
$\mathbf{H}^{(l)}=\left(\mathbf{I}+\alpha\mathbf{A}\right)^{l}\mathbf{H}^{(0)}.$ (5)
@@
$\left(\mathbf{I}+\alpha\mathbf{A}\right)^{l}=\sum_{k}(1+\alpha\lambda_{k})^{l}\mathbf{v_{k}}\mathbf{v_{k}}^{T},$ (6)
@@
${\rm rank}\left(\mathbf{H}^{(l)}\right)\leq{\rm min}\left({\rm rank}\left((\mathbf{I}+\alpha\mathbf{A})^{l}\right),{\rm rank}\left(\mathbf{H}^{(0)}\right)\right)=1,$ (7)
showing that all nodes have the same embedding representations. Considering that the gradient vanishing hinders the convergence of Equation (5), the representations of distinct nodes would not be completely the same on the log loss.
In Figure 2 (a) and (c), we evaluate LightGCN and MF on how the embeddings completely collapse on Yelp when only considering the observed interactions, the parameters are initialized with Xavier initialization. Furthermore, we adopt a Euclidean distance loss implemented on MF without the gradient vanishing issue. We have the following observations:
@@
  • The erank monotonically decreases on all models, and the accuracy barely changes throughout the training, indicating that the embeddings collapse as the training starts and the issue becomes more serious as training proceeds.
@@
    $\mathbf{H}^{(l)}=\left(\mathbf{I}+\alpha\mathbf{A}\right)^{l}\sum_{k=0}^{K}\mathbf{A}^{k}\mathbf{E}.$ (8)
    Here, $\mathbf{E}$ is the initial stacked user/item embeddings sent to the encoder; we ignore the difference between the adjacency matrix used in Equation (5) and LightGCN for simplicity. We can see that increasing the layer $K$ causes the loss function to converge faster, indicating that the over-smoothing in GNN [19] aggravates the collapse issue.
  • The erank tends to converge to a value larger than 1. Besides the gradient vanishing mentioned above, Proposition 1 is based on the assumption that $\mathcal{G}$ is connected, which does not always hold on the real-world datasets. In other words, disconnected nodes (i.e., no reachable paths between them) do not end up having the same representations.
  • By comparing the Euclidean distance loss (MF) and log loss (MF), we can see the erank of Euclidean distance loss with fixed magnitude of gradients whose accuracy is also worse drops more quickly than that of log loss, which further verifies our analysis.
Figure 3: Some analysis on CiteULike. (a) the normalized eigenvalue distribution (Top 500) of $\mathbf{A}-\mathbf{\bar{A}}$; (b) how erank changes on Euclidean distance loss optimizing only positive and both positive and negative samples; (c) how the average scores of sampled positive negative pairs change with training; (d) how erank and recall@10 change with varying negative sampling ratios.
III-B Incomplete Collapse
Existing recommendation algorithms mostly exploit unobserved interactions for optimization. Naturally, we raise a question: Can existing pair-wise learning paradigm completely prevent the collapse? Similarly, we evaluate LightGCN (with BPR), MF (with BPR), and MF (with BCE) on Yelp, and show how erank and accuracy change as training proceeds in Figure 2 (b) and (d). The erank plunges at first, showing a trend similar to Figure 2 (a). Gradually, the erank increases and tends to converge to a value lower than $d$, indicating that the embeddings are negligibly distributed over some dimensions. LightGCN drops more rapidly than MF and converges to a smaller value. In the meanwhile, the accuracy also increases accordingly, and shows a trend similar to erank. It is obvious that existing algorithms (BCE and BPR loss) can alleviate the collapse issue but still cannot completely prevent it. We attempt to analyze how they alleviate the collapse issue and the weakness. Take the BCE loss as an example, the parameters are updated through SGD as follows:
$\mathbf{H}^{(l+1)}=\mathbf{H}^{(l)}+\alpha\left(\mathbf{A}-\mathbf{\bar{A}}\right)\mathbf{H}^{(l)},$ (9)

### v2

@@
Abstract.
Modern recommender systems heavily rely on high-quality representations learned from high-dimensional sparse data. While significant efforts have been invested in designing powerful algorithms for extracting user preferences, the factors contributing to good representations have remained relatively unexplored. In this work, we shed light on an issue in the existing pair-wise learning paradigm (i.e., embedding collapse), that the representations tend to span a subspace of the whole embedding space, leading to a suboptimal solution and reducing the model capacity. Specifically, we show that alignment of positive pairs is equivalent to a low pass filter causing users/items to collapse to a constant vector. While negative sampling can partially mitigate this issue by acting as a high pass filter to balance the spectrum, leading to an incomplete collapse.
To tackle this issue, we present a novel learning paradigm DirectSpec, which directly optimizes the spectrum distribution to ensure that users/items effectively span the entire embedding space. We demonstrate that many self-supervised learning algorithms without explicit negative sampling can be considered as special cases of DirectSpec. Furthermore, we show that optimizing the spectrum inappropriately could also be poisonous to data representation, where the key lies in a dynamic balance between alignment of positive pairs and spectrum balancing. Finally, we propose an enhanced and practical implementation DirectSpec+ to balance the embedding spectrum more adaptively and effectively. We implement DirectSpec+ on two popular recommender models: MF and LightGCN. Our experimental results demonstrate its effectiveness and efficiency over competitive baselines.
@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Information systems Recommender systems
1. Introduction
@@
Recommender systems have penetrated into our daily life, we can see them everywhere such as e-commerce (Wang et al., 2018), short-video (Liu et al., 2019), social network (Jiang et al., 2012), and so on. Collaborative filtering (CF), a fundamental technique for recommendation to discover user preference based on the historical data, has attracted much attention in the last decades. The most extensively used CF technique, matrix factorization (MF) (Koren et al., 2009) which represents users and items as low dimensional latent vectors and estimates the rating as the inner product between latent vectors, has been the cornerstone of modern recommender systems. Since MF estimates the rating with a simple linear function, subsequent works mostly focus on designing more powerful and complex algorithms to model non-linear user-item relations, including but not limited to multi-layer perceptron (MLP) (He et al., 2017), attention mechanism (Kang and McAuley, 2018), reinforcement learning (Zheng et al., 2018), transformer (Sun et al., 2019), diffusion model (Wang et al., 2023), graph neural network (GNN) (Ying et al., 2018; Peng et al., 2022b), etc., and have shown tremendous success.
@@
Although different kinds of methods have been proposed, most of them can be considered as MF variants whose goal is to learn low dimensional representations (with dimension $d$) from the high dimensional sparse interaction matrix (with dimension $D\gg d$). Figure 1 illustrates the top 500 normalized singular value distribution of the interaction matrix of CiteULike (see Section 5.1 for data description). We observe that users/items are predominantly distributed along only a few dimensions in the original space while most dimensions barely contribute (i.e., with singular values close to 0) to the representations. Thus, when users and items are mapped into a more compact embedding space, it is expected that redundant dimensions are all removed and each dimension contributes to the user/item representations as equally and uniformly as possible (i.e., the representations make full use of the embedding space).
@@
Existing recommendation methods mostly utilize positive (observed) and negative (unobserved) samples for optimization. Here, a simple yet fundamental question arises: Can user/item representations of existing works make full use of all dimensions? Unfortunately, by analyzing the spectrum of the embedding matrix, we empirically and theoretically show that users/items tend to span a subspace of the whole embedding space (with dimension $d^{\prime}<d$), where the embeddings collapse along all (complete collapse) or certain dimensions (incomplete collapse). Particularly, alignment of positive pairs is equivalent to a low pass filter, where the representations tend to collapse to a constant vector. Negative sampling is commonly used to optimize recommendation algorithms without causing an explicit embedding collapse by pushing away the unobserved user-item pairs, and we show that it is equivalent to a high pass filter decelerating the speed of collapse by balancing the embedding spectrum. However, the effectiveness of negative sampling is only limited to certain loss functions such as Bayesian personalized ranking (BPR) (Rendle et al., 2009) and binary cross-entropy (BCE) loss (He et al., 2017), where collapse over certain dimensions still exists.
@@
Figure 1. An illustration of the data distribution in the original space and expected distribution in the learning embedding space.
@@
In this work, we tackle the embedding collapse issue from a spectral perspective. We observe that the extent of the collapse is closely related to the spectrum distribution of the embedding matrix. Specifically, only a single singular value dominates when the representation completely collapses, whereas the singular values are uniformly distributed when the representations make full use of the embedding space. Inspired by this observation, we propose to directly optimize the spectrum distribution (dubbed DirectSpec) ensuring that all dimensions equally contribute to the representations. We prove that DirectSpec can completely prevent the embedding collapse and provide a practical implementation with a complexity as $\mathcal{O}(B^{2}d)$ where $B$ is the batch size. In addition, we demonstrate that advanced self-supervised learning (SSL) algorithms such as Barlow Twins, LogDet, and spectral contrastive loss without explicit negative sampling can be considered as special cases of DirectSpec. Furthermore, we show that optimizing the spectrum distribution could also be poisonous to data representation, where the key lies in a dynamic balance between alignment of positive pairs and spectrum balance. Based on this observation, we point out the limitations of DirectSpec and propose an enhanced variant DirectSpec+ to balance the embedding spectrum more adaptively and effectively. Finally, we implement DirectSpec and DirectSpec+ on two popular baselines: MF (Koren et al., 2009) and LightGCN (He et al., 2020), and experimental results show that DirectSpec+ improves BPR and LightGCN by up to 52.6% and 42.4% in terms of nDCG@10, respectively. The contributions of this work can be summarized as follows:
@@
  • We theoretically and empirically show that existing recommendation methods suffer from embedding collapse, where the representations tend to fall into a subspace of the whole embedding space, and analyze the mechanisms causing this issue.
@@
  • We propose a novel learning paradigm DirectSpec which directly optimizes the spectrum distribution, ensuring that users/items effectively span the entire embedding space. We empirically and theoretically show that DirectSpec can prevent embedding collapse.
  • We demonstrate that SSL methods such as Barlow Twins, LogDet, and spectral contrastive loss without explicit negative sampling can be considered as special cases of DirectSpec.
@@
  • Extensive experimental results on three datasets demonstrate the efficiency and effectiveness of our proposed methods.
2. Preliminaries
2.1. Pairwise Learning for CF
@@
Given the user set $\mathcal{U}$ and item set $\mathcal{I}$, the interaction matrix is defined as $\mathbf{R}\in\{0,1\}^{\left|\mathcal{U}\right|\times\left|\mathcal{I}\right|}$, the observed interactions are represented as $\mathbf{R}^{+}=\{r_{ui}=1|u\in\mathcal{U},i\in\mathcal{I}\}$. Users and items are first mapped to low dimensional vectors $\mathbf{H}\in\mathbb{R}^{(\left|\mathcal{U}\right|+\left|\mathcal{I}\right|)\times d}$ through an encoder, where the encoder can be as simple as a linear mapping (Koren et al., 2009) or advanced algorithms such as MLPs (He et al., 2017), GNNs (Wang et al., 2019), etc. Let $\mathbf{h}_{u}$ and $\mathbf{h}_{i}$ be the $u$’s and $i$’s representations, respectively. The goal of CF is to predict unobserved interactions $\mathbf{R}^{-}=\{r_{ui}=0|u\in\mathcal{U},i\in\mathcal{I}\}$ estimated as the inner product between the user and item representations: $\hat{r}_{ui}=\mathbf{h}_{u}^{T}\mathbf{h}_{i}$. The model parameters $\Theta$ are optimized through a loss function $\mathcal{L}$ formulated as follows:
(1) ${\rm arg}\mathop{{\rm min}}\limits_{\mathbf{\Theta}}\mathcal{L}(\hat{r}_{ui},r_{ui}).$
@@
The loss function measures the difference between the estimated score and the ground truth. For instance, BPR and BCE loss are extensively used for CF methods:
(2) $\mathcal{L}_{BPR}=\sum_{(u,i)\in{\mathbf{R}}^{+},(u,j)\in{\mathbf{R}}^{-}}-\ln\sigma(\hat{r}_{ui}-\hat{r}_{uj}),$
@@
[…]
@@
(3) $erank(\mathbf{H})=exp\left(H\left(p_{1},\cdots,p_{d}\right)\right),$
@@
Compared with rank, erank takes the singular value distribution into consideration: the more uniform (sharper) of the distribution, the higher (lower) of the erank (Roy and Vetterli, 2007). The embedding matrix contains the least information when there is only one leading non-zero singular values ($\sigma_{2}=\cdots=\sigma_{d}=0$, and $erank(\mathbf{H})=rank(\mathbf{H})=1$), while erank is maximized when each dimension equally contributes to the representations: $\sigma_{1}=\cdots=\sigma_{d}$ ($erank(\mathbf{H})=rank(\mathbf{H})=d$).
2.2. Graph Representation and Graph Filtering
@@
Graph Filtering. Let $11=\lambda_{1}>\lambda_{2}>\cdots>\lambda_{n}=-$ 1 be the eigenvalues of $\mathbf{A}$ where $n=|\mathcal{V}|$. The components closer to $\lambda_{1}$ and $\lambda_{n}$ correspond to a lower and higher frequency, respectively. $\mathcal{F}^{L}(\mathbf{A})$ is a low pass filter if $|\mathcal{F}^{L}(\lambda_{i})|>|\mathcal{F}^{L}(\lambda_{j})|$ for $\lambda_{i}>\lambda_{j}$ and $\mathcal{F}^{H}(\mathbf{A})$ is a high pass filter if $|\mathcal{F}^{H}(\lambda_{i})|>|\mathcal{F}^{H}(\lambda_{j})|$ for $\lambda_{i}<\lambda_{j}$.
Note that $\mathcal{F}^{L}(\mathbf{A})$ is defined as a low pass filter as it emphasizes more on the lower frequencies and similarly for the definition of the high pass filter. Please refer to (Sandryhaila and Moura, 2014) for detailed introduction of graph frequency.
Figure 2. (a) complete collapse, (b) incomplete collapse, (c) how negative sampling can balance the spectrum on Yelp: the normalized eigenvalue distribution (Top 500) of $\mathbf{A}-\mathbf{\bar{A}}$.
3. Embedding Collapse in CF
3.1. Complete Collapse
Let us first consider a generic loss function only optimizing alignment of positive pairs (i.e., minimizing their distance): $\mathcal{L}(\hat{r}_{ui},r_{ui}),(u,i)\in\mathbf{R}^{+}$.
@@ 3.1. Complete Collapse
(4) $\mathbf{H}^{(l+1)}=\mathbf{H}^{(l)}-\alpha\frac{\partial\mathcal{L}}{\partial\mathbf{H}^{(l)}}=\mathbf{H}^{(l)}+\alpha\mathbf{A}\mathbf{H}^{(l)},$
@@
where $\alpha\in(0,1)$, $\mathbf{A}_{ui}=-\frac{\partial\mathcal{L}}{\partial\hat{r}_{ui}}$ for $(u,i)\in\mathbf{R}^{+}$. Ideally, if $\frac{\partial\mathcal{L}}{\partial\hat{r}_{ui}}$ remain unchanged during training such as the Euclidean distance loss: $\mathcal{L}=\sum\left\|\mathbf{h}_{u}-\mathbf{h}_{i}\right\|^{2}_{2}$, then Equation (4) can be rewritten as:
(5) $\mathbf{H}^{(l)}=\left(\mathbf{I}+\alpha\mathbf{A}\right)\mathbf{H}^{(l-1)}=\cdots=\left(\mathbf{I}+\alpha\mathbf{A}\right)^{l}\mathbf{H}^{(0)}.$
@@
(6) $\left(\mathbf{I}+\alpha\mathbf{A}\right)^{l}=\sum_{k}(1+\alpha\lambda_{k})^{l}\mathbf{v_{k}}\mathbf{v_{k}}^{T},$
@@
(7) ${\rm rank}\left(\mathbf{H}^{(l)}\right)\leq{\rm min}\left({\rm rank}\left((\mathbf{I}+\alpha\mathbf{A})^{l}\right),{\rm rank}\left(\mathbf{H}^{(0)}\right)\right)=1,$
[…]
@@
  • The erank monotonically declines on both models, indicating that the embeddings collapse as the training starts and the issue becomes more serious as training proceeds.
@@
    (8) $\mathbf{H}^{(l)}=\left(\mathbf{I}+\alpha\mathbf{A}\right)^{l}\sum_{k=0}^{K}\mathbf{A}^{k}\mathbf{E}.$
    Here, $\mathbf{E}$ is the initial stacked user/item embeddings sent to the encoder; we ignore the difference between the adjacency matrix used in Equation (5) and LightGCN for simplicity. We can see that increasing the layer $K$ causes the loss function to converge faster, indicating that the over-smoothing in GNN (Li et al., 2018) aggravates the collapse issue.
  • The erank tends to converge to a value larger than 1 due to the two reasons we mentioned above. By comparing the Euclidean distance loss (MF) and log loss (MF), we can see the erank of the Euclidean distance loss with stable gradients drops more quickly than that of log loss, which further verifies our analysis.
3.2. Incomplete Collapse
[…]
(9) $\mathbf{H}^{(l+1)}=\mathbf{H}^{(l)}+\alpha\left(\mathbf{A}-\mathbf{\bar{A}}\right)\mathbf{H}^{(l)},$